# Les Trente Glorieuses: 29 years of helioseismic observations with the Luminosity Oscillations Imager

### v1

@@ Abstract
Helioseismology started in the late 1970’s with the first detection of global solar p modes by Claverie et al. (1979), confirmed by Grec et al. (1980), using radial velocity measurements; and by Woodard & Hudson (1983) using irradiance measurements with ACRIM111Active Cavity Radiometer Irradiance Monitor (Willson, 1979) onboard SMM222Solar Maximum Mission. These breakthroughs were followed by the first inversion of the solar structure performed by Gough (1984) and Duvall et al. (1984) using the so-called Duvall’s law (Duvall & Harvey, 1984). Helioseismology really became a useful tool when it demonstrated that the equation of state could be in error, and was indeed in error (Kosovichev et al., 1992).
@@
The diagnostic potential of helioseismology for our physic had already been perceived in the mid 80’s when the space mission SoHO333Solar and Heliospheric Observatory, an ESA/NASA mission (Special volumes of Solar Physics, 1995, 162;1997, 170; et 1997,175) was approved. The need for uninterrupted data can be fulfilled by a space mission, but also by ground-based networks of instrument, networks on which the Sun never sets: GONG444Global Oscillation Network Group (dedicated volume of Science, 1996, 272), IRIS555Investigation on the Rotation and Interior of the Sun (Egamberdiev & Fossat, 1991) and BiSON666Birmingham Solar-Oscillations Network (Elsworth et al., 1991).
@@
The detection of solar oscillations using solar radial velocities is feasible from the ground. It is not so obvious how to perform such measurements using irradiance or radiance fluctuations because of transmission variations through the Earth atmosphere. Following the detection of p modes with the ACRIM instrument, Fröhlich & Van der Raay (1984) used photometers aboard balloons to reduce the impact of the Earth atmosphere on the measurement. A similar photometer (IPHIR777InterPlanetary Helioseismology by IRradiance measurements, (Fröhlich et al., 1988)) was flown aboard the PHOBOS II888Soviet Mars exploration mission, (See Siddiqi, 2016) mission. This instrument returned excellent helioseismic measurements as shown by Toutain & Fröhlich (1992).
@@
In 1987, the SoHO mission was selected by the European Space Agency as part of it Horizon 2 programme. The mission consisted of in-situ, coronal, particle, and helioseismic instruments. Following the at-the-times recent developments in helioseismology, three helioseismic instruments were included for obtaining solar radial-velocity and irradiance measurements using either imagers or looking at the Sun as a star. GOLF999Global Oscillations at Low Frequency (Gabriel et al., 1995) uses a Sodium cell to measure solar radial velocities look at the Sun as a star, while MDI 101010Michelson Doppler Imager (Scherrer et al., 1995) uses a one-million pixel detector for obtaining spatially resolved data. The VIRGO111111Variability of IRradiance and Gravity Oscillations (Fröhlich et al., 1995) instrument measures the irradiance of the Sun as a star, the spectral radiance with photometers, and the spectral radiance spatially resolved over the solar disk. The latter is performed by the Luminosity Oscillations Imager (LOI) whose concept was described by Andersen et al. (1988), and later built as a space instrument (Appourchaux et al., 1997).
@@
The LOI is one of a kind, that is the only instrument that detected in intensity the low degree p modes from the ground (Appourchaux et al., 1995), and that has been measuring the low degree p modes from space since 1996. Its design was specifically aiming at detecting modes with a different sensitivity for degree greater than 1; thereby optimizing the chance of detecting the solar g modes using different signals. Instruments that detects p modes in velocity, such as GOLF and BiSON (Broomhall et al., 2007) or HMI121212Helioseismic and Magnetic Imager (Scherrer et al., 2012) and GONG (Appourchaux, 2024), benefit from a better signal-to-noise ratio allowing them to detect modes below 1 Hz. In addition, instruments such as GONG, MDI and HMI make images of the Sun permitting them to derive by inversion the internal structure from the surface below the convection zone. Nevertheless, it must be emphasized that scientific progress should not always be ranked on the ability of making interesting discoveries, but also on the ability of confirming new discoveries. The failure to confirm g-mode detection can serve as an example (Appourchaux et al., 2000, 2010).
@@ 3.1 Roll maneuver correction
After May 25, 2003, the HGA failed to move for sending data towards the Earth. It was then decided to leave the HGA in a fixed position and rotate the spacecraft by 180 degrees every 3 months for transmitting to Earth131313See https://soho.nascom.nasa.gov/soc/soho_events/SOHO-Spacecraft-Events.csv. When the S/C is rolled by 180 degrees, the pixel numbering is redefined in the process, i.e. pixels 1, 2, 5, 6, 7, 8 becoming pixels 4, 3, 8, 7, 6, 5 and vice versa; and the guiding pixels 13, 14 becoming 16, 15 and vice versa.
@@ 4.2 Spherical harmonics extraction
For combining the pixel data, I used the spherical harmonic decomposition computed as in Appourchaux & Andersen (1990), using the limb darkening from Allen (1975). The filters are normalized according to Eq. (17) of Appourchaux et al. (1998a). This normalization allows a symmetrical leakage matrix (Appourchaux et al., 1998a). Example for such leakage matrix can be found in Appourchaux et al. (1998c). Since the apparent solar diameter changes along the orbit, the spherical harmonic filters vary slowly with time. To alleviate this problem, time-independent filters are applied to the pixels for extracting the modes. The filters are computed as an average over one year of filters computed weekly. Over this period the angle has a mean close to zero. Gizon et al. (1998) showed that it was possible to detect the influence of a non-zero angle on the p-mode data. Therefore, this average minimizes the effect of the angle on the data. After combining the pixels with the filters, the time series are Fourier transformed; the positive frequencies provide the signal for , and the negative frequencies provide the conjugate of the signal for Appourchaux et al. (1998a).
@@ 5.1 Power spectra fitting
The code used by Appourchaux et al. (2008) for the CoRoT star HD49933 was indeed an IDL code. There were already attempts to write a C++ code (progFIT) based on this IDL code (Neiner & Appourchaux, 2004) which was not yet available in 2008. The development of that code continued until 2010, but the code was not still fully functioning (Neiner, private communication). The code was then picked up from its ashes in 2021 when its use for the Planetary transits and oscillations of stars (PLATO) mission was considered an asset given the large sample of solar-like stars (About more than 20, stars for the core P1 sample, Rauer et al., 2025). The code was then made fully functional and was tested for the PLATO mission on hundreds of solar-like stars. The code can be found on Gitlab141414git.ias.u-psud.fr/plato_pdc/plato_wp37/wp372_msap3
@@
Following the successful development of progFIT, another code was developed for fitting the NS and EW difference signals, and HV guiding signals. Since these signals are by construction not sensitive to =0 modes, the new code (guiFIT) took into account this lack of visibility, but also the different visibilities in a multiplet. The code can also be found on Gitlab151515git.ias.u-psud.fr/plato_pdc/plato_wp37/wp372_msap3
@@ 5.2 Fourier spectrum fitting
The 10.7-cm flux could be downloaded at the SIDC (Solar Influence Data Analysis Center) in 2023 with a time series running from April 1st 1996 to March 31st, 2021; the data are at the time of writing no longer available. The flux can now be found daily thanks to Space Weather Canada161616spaceweather.gc.ca/forecast-prevision/solar-solaire/solarflux/sx-5-en.php, and the two time series were merged into a single time series from April 1st 1996 to March 31st, 2025, for producing a yearly average. Since I used the 29 yearly time series of the different signals for that study, I also averaged the flux over the same period of one year. The 29 power spectra were fitted using a variable and a fixed noise background. The two fits were used to reduce the bias on linewidths and mode heights that occurs when one uses a different background (See for instance Appourchaux et al., 2014).
@@ 9.1 Limb darkening and solar radius calibration
The computation of the mode visibilities requires calibrating the limb darkening and the effective size of the solar radius on the LOI detector (See Appendix B). The calibration was done by computing the ratio of the mean of the NS pixels (1, 2. 3, 4) to the Central pixels (9, 10, 11, 12), the ratio of the mean of the EW pixels (5, 6, 7, 8) to the Central pixels, and the ratio of the mean of the guiding pixels (12, 13, 14, 15) as a function of the SoHO-Sun distance in AU. The computation of the relative fluxes was performed using the shape of the pixels and the presence of the tracks as provided by the manufacturer171717Advanced Micro Electronics, Horten, Norway. The tracks are non sensitive part of the detector required for getting the current resulting from the photoelectric effect. The limb darkening is assumed to be expressed as:
@@ 10 Discussion and conclusion
I also studied the heights of the modes detected in the guiding pixels and HV signals. The mode heights in the HV guiding signals are about 5-10 times larger than the modes observed in FD for , while it is about 10-20s times larger for (after correction of the FD visibility). Such higher mode heights are in agreement with other observations performed by Toner et al. (1999) and Corbard et al. (2013) with MDI, HMI and the PICARD mission. On the other hand, the visibility ratios for the EW and NS signals (guiding pixels and HV) reported in this article are clearly not in agreement with the expectation of the simpleI theory provided in Appendix C. There are several discrepancies with theory to be noted (See Table D.4 and D.5): 1) the mode visibilities of the EW signals are systematically higher than those of the NS signals; 2) the mode visibilities of the modes are significantly higher than those of the modes for the guiding pixels and the HV signals. If one compares the results of the HV signals with the EW and NS difference (See Table D;3), it is clear that for these latter signals 1) the mode visibilities in EW and NS are identical for and ; and 2) the ratio of mode visibilities of the to the modes are in agreement with theory. It means that the intensity perturbation at the center of the disk () down to follows a product of the limb darkening with a spherical harmonic (Eq. C.1).
When this is not the case, one can use the formalism given by Toutain et al. (1999) for deriving the visibility ratio of with respect to . I show in Appendix C that the ratio of these visibilities for either the EW or NS signals is . For spherical harmonics, this ratio is about 0.2 (See Table D.4). Using the relative intensity perturbation provided by Toutain et al. (1999) in their Figures 5 and 6 for (used by ) and for (used by ), one can deduce that the ratio should be much less than 1 as well. The intensity perturbations provided by Toutain et al. (1999) do not contradict the simple theory provided by Table D.4. The theory of Toutain et al. (1999) was revised by Prokhorov et al. (2014) and Kostogryz et al. (2021). The latter used additional opacity in addition to the sole contribution of bound-free transitions of H- included in the model of Toutain et al. (1999). Unfortunately, Kostogryz et al. (2021) did not provide the intensity perturbations for and . The computation of such perturbation using the methodology of Kostogryz et al. (2021) is beyond the scope of the current article.
@@ Acknowledgements.
- Fournier et al. (2025) Fournier, D., Kostogryz, N. M., Gizon, L., et al. 2025, arXiv e-prints, arXiv:2507.13756
@@ C.7 The multiplet
In this case, the responses for the EW / NS pixels and the HV EW signals are different from zero, while nothing is transmitted for . For the NS HV signal, only is transmitted. The visibilities111sum over the transmitted modes in the power spectrum for the EW signals and for the NS signals are identical.

### v2

@@ Abstract
Helioseismology started in the late 1970’s with the first detection of global solar p modes by Claverie et al. (1979), confirmed by Grec et al. (1980), using radial velocity measurements; and by Woodard & Hudson (1983) using irradiance measurements with ACRIM Monitor (Active Cavity Radiometer Irradiance, Willson, 1979) onboard SMM (Solar Maximum Mission). These breakthroughs were followed by the first inversion of the solar structure performed by Gough (1984) and Duvall et al. (1984) using the so-called Duvall’s law (Duvall & Harvey, 1984). Helioseismology really became a useful tool when it demonstrated that the equation of state could be in error, and was indeed in error (Kosovichev et al., 1992).
@@
The diagnostic potential of helioseismology for our physic had already been perceived in the mid 80’s when the space mission SoHO (Solar and Heliospheric Observatory, an ESA/NASA mission, Special volumes of Solar Physics, 1995, vol. 162;1997, vol. 170; and 1997,vol. 175)) was approved. The need for uninterrupted data can be fulfilled by a space mission, but also by ground-based networks of instrument, networks on which the Sun never sets: GONG (Global Oscillation Network Group, dedicated volume of Science, 1996, vol. 272), IRIS (Investigation on the Rotation and Interior of the Sun, Egamberdiev & Fossat, 1991) and BiSON (Birmingham Solar-Oscillations Network, Elsworth et al., 1991).
@@
The detection of solar oscillations using solar radial velocities is feasible from the ground. It is not so obvious how to perform such measurements using irradiance or radiance fluctuations because of transmission variations through the Earth atmosphere. Following the detection of p modes with the ACRIM instrument, Fröhlich & Van der Raay (1984) used photometers aboard balloons to reduce the impact of the Earth atmosphere on the measurement. A similar photometer (IPHIR: InterPlanetary Helioseismology by IRradiance measurements, Fröhlich et al., 1988) was flown aboard PHOBOS II (Soviet Mars exploration mission, see Siddiqi, 2016). This instrument returned excellent helioseismic measurements as shown by Toutain & Fröhlich (1992).
@@
In 1987, the SoHO mission was selected by the European Space Agency as part of it Horizon 2 programme. The mission consisted of in-situ, coronal, particle, and helioseismic instruments. Following the at-the-times recent developments in helioseismology, three helioseismic instruments were included for obtaining solar radial-velocity and irradiance measurements using either imagers or looking at the Sun as a star. GOLF (Global Oscillations at Low Frequency, Gabriel et al., 1995) uses a Sodium cell to measure solar radial velocities look at the Sun as a star, while MDI (Michelson Doppler Imager, Scherrer et al., 1995) uses a one-million pixel detector for obtaining spatially resolved data. The VIRGO (Variability of IRradiance and Gravity Oscillations, Fröhlich et al., 1995) instrument measures the irradiance of the Sun as a star, the spectral radiance with photometers, and the spectral radiance spatially resolved over the solar disk. The latter is performed by the Luminosity Oscillations Imager (LOI) whose concept was described by Andersen et al. (1988), and later built as a space instrument (Appourchaux et al., 1997).
@@
The LOI is one of a kind, that is the only instrument that detected in intensity the low degree p modes from the ground (Appourchaux et al., 1995), and that has been measuring the low degree p modes from space since 1996. Its design was specifically aiming at detecting modes with a different sensitivity for degree greater than 1; thereby optimizing the chance of detecting the solar g modes using different signals. Instruments that detects p modes in velocity, such as GOLF and BiSON (Broomhall et al., 2007) or HMI (Helioseismic and Magnetic Imager, Scherrer et al., 2012) and GONG (Appourchaux, 2024), benefit from a better signal-to-noise ratio allowing them to detect modes below 1 Hz. In addition, instruments such as GONG, MDI and HMI make images of the Sun permitting them to derive by inversion the internal structure from the surface below the convection zone. Nevertheless, it must be emphasized that scientific progress should not always be ranked on the ability of making interesting discoveries, but also on the ability of confirming new discoveries. The failure to confirm g-mode detection can serve as an example (Appourchaux et al., 2000, 2010).
@@ 3.1 Roll maneuver correction
After May 25, 2003, the HGA failed to move for sending data towards the Earth. It was then decided to leave the HGA in a fixed position and rotate the spacecraft by 180 degrees every 3 months for transmitting to Earth111See https://soho.nascom.nasa.gov/soc/soho_events/SOHO-Spacecraft-Events.csv. When the S/C is rolled by 180 degrees, the pixel numbering is redefined in the process, i.e. pixels 1, 2, 5, 6, 7, 8 becoming pixels 4, 3, 8, 7, 6, 5 and vice versa; and the guiding pixels 13, 14 becoming 16, 15 and vice versa.
@@ 4.2 Spherical harmonics extraction
For combining the pixel data, I used the spherical harmonic decomposition computed as in Appourchaux & Andersen (1990), using the limb darkening from Allen (1975). The filters are normalized according to Eq. (17) of Appourchaux et al. (1998a). This normalization allows a symmetrical leakage matrix (Appourchaux et al., 1998a). Example for such leakage matrix can be found in Appourchaux et al. (1998c). Since the apparent solar diameter changes along the orbit, the spherical harmonic filters vary slowly with time. To alleviate this problem, time-independent filters are applied to the pixels for extracting the modes. The filters are computed as an average over one year of filters computed weekly. Over this period the angle has a mean close to zero. Gizon et al. (1998) showed that it was possible to detect the influence of a non-zero angle on the p-mode data. Therefore, this average minimizes the effect of the angle on the data. After combining the pixels with the filters, the time series are Fourier transformed; the positive frequencies provide the signal for , and the negative frequencies provide the conjugate of the signal for (Appourchaux et al., 1998a).
@@ 5.1 Power spectra fitting
The code used by Appourchaux et al. (2008) for the CoRoT star HD49933 was indeed an IDL code. There were already attempts to write a C++ code (progFIT) based on this IDL code (Neiner & Appourchaux, 2004) which was not yet available in 2008. The development of that code continued until 2010, but the code was not still fully functioning (Neiner, private communication). The code was then picked up from its ashes in 2021 when its use for the Planetary transits and oscillations of stars (PLATO) mission was considered an asset given the large sample of solar-like stars (About more than 20, stars for the core P1 sample, Rauer et al., 2025). The code was then made fully functional and was tested for the PLATO mission on hundreds of solar-like stars. The code can be found on Gitlab222git.ias.u-psud.fr/plato_pdc/plato_wp37/wp372_msap3
@@
Following the successful development of progFIT, another code was developed for fitting the NS and EW difference signals, and HV guiding signals. Since these signals are by construction not sensitive to =0 modes, the new code (guiFIT) took into account this lack of visibility, but also the different visibilities in a multiplet. The code can also be found on Gitlab333git.ias.u-psud.fr/plato_pdc/plato_wp37/wp372_msap3
@@ 5.2 Fourier spectrum fitting
The 10.7-cm flux could be downloaded at the SIDC (Solar Influence Data Analysis Center) in 2023 with a time series running from April 1st 1996 to March 31st, 2021; the data are at the time of writing no longer available. The flux can now be found daily thanks to Space Weather Canada444spaceweather.gc.ca/forecast-prevision/solar-solaire/solarflux/sx-5-en.php, and the two time series were merged into a single time series from April 1st 1996 to March 31st, 2025, for producing a yearly average. Since I used the 29 yearly time series of the different signals for that study, I also averaged the flux over the same period of one year. The 29 power spectra were fitted using a variable and a fixed noise background. The two fits were used to reduce the bias on linewidths and mode heights that occurs when one uses a different background (See for instance Appourchaux et al., 2014).
@@ 9.1 Limb darkening and solar radius calibration
The computation of the mode visibilities requires calibrating the limb darkening and the effective size of the solar radius on the LOI detector (See Appendix refAppendixB). The calibration was done by computing the ratio of the mean of the NS pixels (1, 2. 3, 4) to the Central pixels (9, 10, 11, 12), the ratio of the mean of the EW pixels (5, 6, 7, 8) to the Central pixels, and the ratio of the mean of the guiding pixels (12, 13, 14, 15) as a function of the SoHO-Sun distance in AU. The computation of the relative fluxes was performed using the shape of the pixels and the presence of the tracks as provided by the manufacturer555Advanced Micro Electronics, Horten, Norway. The tracks are non sensitive part of the detector required for getting the current resulting from the photoelectric effect. The limb darkening is assumed to be expressed as:
@@ 10 Discussion and conclusion
I also studied the heights of the modes detected in the guiding pixels and HV signals. The mode heights in the HV guiding signals are about 5-10 times larger than the modes observed in FD for , while it is about 10-20s times larger for (after correction of the FD visibility). Such higher mode heights are in agreement with other observations performed by Toner et al. (1999) and Corbard et al. (2013) with MDI, HMI and the PICARD mission. On the other hand, the visibility ratios for the EW and NS signals (guiding pixels and HV) reported in this article are clearly not in agreement with the expectation of the simple theory provided in Appendix C. There are several discrepancies with theory to be noted (See Table 14 and 15): 1) the mode visibilities of the EW signals are systematically higher than those of the NS signals; 2) the mode visibilities of the modes are significantly higher than those of the modes for the guiding pixels and the HV signals. If one compares the results of the HV signals with the EW and NS difference (See Table 13), it is clear that for these latter signals 1) the mode visibilities in EW and NS are identical for and ; and 2) the ratio of mode visibilities of the to the modes are in agreement with theory. It means that the intensity perturbation at the center of the disk () down to follows a product of the limb darkening with a spherical harmonic (Eq. C.1).
When this is not the case, one can use the formalism given by Toutain et al. (1999) for deriving the visibility ratio of with respect to . I show in Appendix C that the ratio of these visibilities for either the EW or NS signals is . For spherical harmonics perturbations, this ratio is about 0.2 (See Table 14). Using the relative intensity perturbation provided by Toutain et al. (1999) in their Figures 5 and 6 for (used by ) and for (used by ), one can deduce that the ratio should be much less than 1 as well. The intensity perturbations provided by Toutain et al. (1999) do not contradict the simple theory provided by Table 14. The theory of Toutain et al. (1999) was revised by Prokhorov et al. (2014) and Kostogryz et al. (2021). The latter used additional opacity in addition to the sole contribution of bound-free transitions of H- included in the model of Toutain et al. (1999). Unfortunately, Kostogryz et al. (2021) did not provide the intensity perturbations for and . The computation of such perturbation using the methodology of Kostogryz et al. (2021) is beyond the scope of the current article.
@@ Acknowledgements.
- Fournier et al. (2025) Fournier, D., Kostogryz, N., Gizon, L., et al. 2025, A&A, 702, A253
@@ C.7 The multiplet
In this case, the responses for the EW / NS pixels and the HV EW signals are different from zero, while nothing is transmitted for . For the NS HV signal, only is transmitted. The visibilities666sum over the transmitted modes in the power spectrum for the EW signals and for the NS signals are identical.